Prove that

(i) the three points A(1, 4, 2), B(3, 2, 4) and C(5, 0, 6) are collinear,

(ii) the four points P(2, ;1, 1), Q(1, 3, ;2), R((2, 1, ;3) and S(3, 2, 0) are coplanar


1
Expert's answer
2021-02-02T01:26:32-0500

"\left(i\right)\:If\:three\:direction\:ratios\:of\:AB\:and\:BC\:are\:proportional,"

"then\:the\:points\:A,\:B\:and\:C\:are\:collinear."

"Direction\:ratios\:are:\:"

"\vec{AB}\:=B-A"

"=\left(3-1\right),\left(2-4\right),\left(4-2\right)"

"=2,-2,2"

"so\:a_1=2\:\:b_1=-2\:c_1=2"

"\vec{BC}=C-B"

"=\left(5-3\right),\left(0-2\right),\left(6-4\right)"

"=2,-2,2"

"so\:a_2=2,\:b_2=-2\:c_2=2"

"\frac{a_2}{a_1}=\frac{2}{2}=1\:"

"\frac{b_2}{b_1}=\frac{-2}{-2}=1"

"\frac{c_2}{c_1}=\frac{2}{2}=1"

"since\:\frac{a_2}{a_1}=\frac{b_2}{b_1}=\frac{c_2}{c_1}=1"

"Therefore\:A,B\:and\:C\:are\:collinear."

"\left(ii\right)Given\:that\:points\:A\left(2,2,1\right),\:Q\left(1,3,2\right),\:R\left(2,1,3\right)and\:S\left(3,2,0\right)"

"\vec{AQ}=Q-A=\left(-i+2j+k\right)"

"\vec{AR}=R-A=\left(2k\right)"

"\vec{AS}=S-A=\left(i+j-k\right)"

"Therefore\:\left|\vec{AQ}\:\vec{AR}\:\vec{AS}\right|=\begin{pmatrix}-1&2&1\\ 0&0&2\\ 1&1&-1\end{pmatrix}"

"=\:-1\left(0-2\right)+2\left(2-0\right)+1\left(0-0\right)"

"=2+4+0=6\:"

"\:Hence,\:the\:points\:are\:not\:coplanar."